# Optimal Quantization on Spherical Surfaces: Continuous and Discrete Models – A Beginner-Friendly Expository Study

### v1

@@ 3. Quantization Framework
Assume $M$ is a smooth curve with arc-length parameter $s$ and geodesic distance $d_{G}$, and $P$ has a continuous density $\rho$ with respect to $ds$. If $\alpha^{\ast}=\{a_{i}^{\ast}\}$ is optimal and $R_{i}$ is the Voronoi region of $a_{i}^{\ast}$, then $a_{i}^{\ast}$ minimizes
$F_{i}(a)\;=\;\int_{R_{i}}d_{G}(x,a)^{2}\,\rho(x)\,ds$
with respect to $a$ lying on the curve. In particular, when $R_{i}$ is a geodesic interval and $d_{G}$ coincides with interval length, $a_{i}^{\ast}$ is the midpoint of $R_{i}$ when $\rho$ is constant.
@@
$\int_{s_{L}}^{s_{R}}(s-s_{0})\,\rho(s)\,ds=0,$
i.e., $s_{0}$ is the intrinsic (conditional) mean of $s$ over the cell. In particular, if $\rho$ is constant, symmetry forces $s_{0}=(s_{L}+s_{R})/2$, so $a$ is the midpoint of $R_{i}$. The second variation satisfies
$\frac{d^{2}}{ds_{0}^{2}}F_{i}(a)=2\int_{s_{L}}^{s_{R}}\rho(s)\,ds>0,$
@@
Remark $3.$
Equation (6) states: compute the weighted Euclidean mean of the assigned points and project it back to the sphere by normalization. If $P$ is a continuous Borel probability measure, then the result follows analogously, see (5).
@@ 4.2. Optimality of uniform partitions
Theorem 4.3 (Equator: structure of optimal $n$-means).
@@ 4.4. Exact quantization error on the equator
Theorem 4.5 (Equator: explicit $V_{n}$).
@@
By Theorem 4.3, the optimal partition has $n$ congruent arcs of length $h=L/n=2\pi/n$, each represented by its midpoint. On one cell, let $t\in[-h/2,h/2]$ denote the geodesic coordinate relative to the midpoint. The conditional mean squared error on that cell equals
@@
Remark 4.6 (Relation to the line).
@@ 5.1. Geometry and effective metric
Theorem 5.2 (Small circle: structure and error).
@@
The proof is identical to the equator case, replacing the total length $L$ by $L_{\lambda}=2\pi\cos\lambda$. Rotation invariance around the axis through the poles yields equally spaced codepoints; uniform density and geodesic intervals force midpoints as representatives. The one-cell computation gives $h=L_{\lambda}/n$, and the same calculation as in Theorem 4.5 yields $V_{n}=L_{\lambda}^{2}/(12n^{2})$. ∎
Remark 5.3 (Latitude effect).
@@ 6.1. Uniform continuous model on an arc
Theorem 6.2 (Uniform arc: structure and error).
@@
As in Theorem 4.3, convexity of the squared-distance on geodesic intervals and the uniform density imply that each Voronoi cell must be a contiguous interval and the representative is its midpoint (Proposition 3.1). If two adjacent cells had unequal lengths, transferring an $\varepsilon$-slice from the larger cell to the smaller one strictly decreases the total cost by a standard “balancing” calculation, contradicting optimality. Hence all cells have length $h=L/n$. The one-cell calculation with midpoints yields the error $h^{2}/12$ per cell in conditional mean, and aggregating over $n$ cells gives $L^{2}/(12n^{2})$ as before. ∎
Remark 6.3 (From arc to full circle).
Taking $L\to 2\pi$ recovers Theorem 4.5. The only difference between the arc and circle is the wrap-around periodicity; the per-cell analysis is identical.
@@ 6.4. Finite discrete uniform model on an arc
(i) By convexity of the one-block objective and the exchange argument from Theorem 6.2, blocks must be contiguous and as balanced as possible. (ii) With uniform spacing $\Delta=L/(m-1)$ (or approximately $L/m$ for large $m$), the sum of squared distances in a block centered at its midpoint is a symmetric discrete quadratic sum. Dividing by $m$ and summing over $n$ blocks yields the stated formula, which approaches the continuous integral $h^{2}/12$ per block with $h=L/n$ as $m\to\infty$. ∎
@@ 7.1. A one-line principle
Proposition 7.2 (Uniform one-dimensional geodesic principle).
@@
7.3. Summary table
@@
Remark 7.4 (Chordal vs. geodesic distance).
@@ 8. Worked examples
Example 8.1 (Equator with $n=3$).
@@
Example 8.5 (Antipodal pair).
Let $x_{1}=(1,0,0)$, $x_{2}=(-1,0,0)$ with $p_{1}=p_{2}=\tfrac{1}{2}$. For $n=1$, the optimal point $a^{*}$ lies at $(0,1,0)$ or (0,1,0)(0,-1,0), giving $d_{G}(x_{i},a^{*})=\pi/2$ and
$V_{1}(P)=\frac{\pi^{2}}{4}.$
For $n=2$, $\alpha=\{x_{1},x_{2}\}$ yields $V_{2}(P)=0$.
@@ 11. Further Remarks and Extensions
  • Nonuniform densities. For nonuniform distributions on an arc or circle, the Voronoi cells are no longer equal in length; their boundaries shift to equalize weighted distortion. The analysis involves solving $\rho(x)(x-a_{i})$ equilibrium equations along geodesics.

### v2

@@ 3. Quantization Framework
Assume $M$ is a smooth curve with arc-length parameter $s$ and geodesic distance $d_{G}$, and $P$ has a continuous density $f$ with respect to $ds$. If $\alpha^{\ast}=\{a_{i}^{\ast}\}$ is optimal and $R_{i}$ is the Voronoi region of $a_{i}^{\ast}$, then $a_{i}^{\ast}$ minimizes
$F_{i}(a)\;=\;\int_{R_{i}}d_{G}(x,a)^{2}\,f(x)\,ds$
with respect to $a$ lying on the curve. In particular, when $R_{i}$ is a geodesic interval and $d_{G}$ coincides with interval length, $a_{i}^{\ast}$ is the midpoint of $R_{i}$ when $f$ is constant.
@@
$\int_{s_{L}}^{s_{R}}(s-s_{0})\,f(s)\,ds=0,$
i.e., $s_{0}$ is the intrinsic (conditional) mean of $s$ over the cell. In particular, if $f$ is constant, symmetry forces $s_{0}=(s_{L}+s_{R})/2$, so $a$ is the midpoint of $R_{i}$. The second variation satisfies
$\frac{d^{2}}{ds_{0}^{2}}F_{i}(a)=2\int_{s_{L}}^{s_{R}}f(s)\,ds>0,$
@@
Remark 3.3.
Equation (7) states: compute the weighted Euclidean mean of the assigned points and project it back to the sphere by normalization. If $P$ is a continuous Borel probability measure, then the result follows analogously, see (5). This is the discrete analogue of the extrinsic centroid condition discussed in Section 1.13.
@@ 4.2. Optimality of uniform partitions
Theorem 4.2.1 (Equator: structure of optimal $n$-means).
@@
Remark 4.2.2.
(Relation to classical one-dimensional quantization). The argument above is an intrinsic reformulation of the classical one-dimensional uniform quantization problem. Under arc-length parameterization, the equator $\Gamma$ is isometric to a circle of length $L=2\pi$, and the squared geodesic distance coincides with the squared Euclidean distance on this circle modulo periodicity. It is a standard result in one-dimensional Euclidean quantization theory that, for a uniform distribution on an interval or circle of length $L$, the optimal $n$-means are equally spaced points and the quantization error equals $V_{n}=L^{2}/(12n^{2})$; see, for example, Graf and Luschgy [4]. This identification completes the argument on $\Gamma$ in a fully rigorous manner.
@@ 4.3. Exact quantization error on the equator
Theorem 4.3.1 (Equator: explicit $V_{n}$).
@@
By Theorem 4.2.1, the optimal partition has $n$ congruent arcs of length $h=L/n=2\pi/n$, each represented by its midpoint. On one cell, let $t\in[-h/2,h/2]$ denote the geodesic coordinate relative to the midpoint. The conditional mean squared error on that cell equals
@@
Remark 4.3.2 (Relation to the line).
@@ 5.1. Geometry and effective metric
Theorem 5.1.1 (Small circle: structure and error).
@@
The proof is identical to the equator case, replacing the total length $L$ by $L_{\lambda}=2\pi\cos\lambda$. Rotation invariance around the axis through the poles yields equally spaced codepoints; uniform density and geodesic intervals force midpoints as representatives. The one-cell computation gives $h=L_{\lambda}/n$, and the same calculation as in Theorem 4.3.1 yields $V_{n}=L_{\lambda}^{2}/(12n^{2})$. ∎
Remark 5.1.2 (Latitude effect).
@@ 6.1. Uniform continuous model on an arc
Theorem 6.1.1 (Uniform arc: structure and error).
@@
As in Theorem 4.2.1, convexity of the squared-distance on geodesic intervals and the uniform density imply that each Voronoi cell must be a contiguous interval and the representative is its midpoint (Proposition 3.1). If two adjacent cells had unequal lengths, transferring an $\varepsilon$-slice from the larger cell to the smaller one strictly decreases the total cost by a standard “balancing” calculation, contradicting optimality. Hence all cells have length $h=L/n$. The one-cell calculation with midpoints yields the error $h^{2}/12$ per cell in conditional mean, and aggregating over $n$ cells gives $L^{2}/(12n^{2})$ as before. ∎
Remark 6.1.2 (From arc to full circle).
Taking $L\to 2\pi$ recovers Theorem 4.3.1. The only difference between the arc and circle is the wrap-around periodicity; the per-cell analysis is identical.
@@ 6.2. Finite discrete uniform model on an arc
(i) By convexity of the one-block objective and the exchange argument from Theorem 6.1.1, blocks must be contiguous and as balanced as possible. (ii) With uniform spacing $\Delta=L/(m-1)$ (or approximately $L/m$ for large $m$), the sum of squared distances in a block centered at its midpoint is a symmetric discrete quadratic sum. Dividing by $m$ and summing over $n$ blocks yields the stated formula, which approaches the continuous integral $h^{2}/12$ per block with $h=L/n$ as $m\to\infty$. ∎
@@ 7.1. A one-line principle
Proposition 7.1.1 (Uniform one-dimensional geodesic principle).
@@
The unified reduction principle described above is illustrated explicitly in the summary presented in Table 1.
7.2. Table 1
@@
Remark 7.2.1 (Chordal vs. geodesic distance).
@@ 8. Worked examples
Example 8.1 (Equator $\Gamma$ with $n=3$).
@@
Example 8.5 (Antipodal pair: intrinsic vs. extrinsic centroids).
Let
$x_{1}=(1,0,0),\qquad x_{2}=(-1,0,0)$
with equal weights $p_{1}=p_{2}=\tfrac{1}{2}$, and let
$P=\tfrac{1}{2}(\delta_{x_{1}}+\delta_{x_{2}})$
be a discrete probability measure on $\mathbb{S}^{2}$.
Intrinsic (Fréchet) formulation. For $n=1$, we minimize the Fréchet functional
$F(a)=\tfrac{1}{2}d_{G}(x_{1},a)^{2}+\tfrac{1}{2}d_{G}(x_{2},a)^{2},\qquad a\in\mathbb{S}^{2}.$
Since $x_{1}$ and $x_{2}$ are antipodal, for every $a\in\mathbb{S}^{2}$,
$d_{G}(x_{1},a)+d_{G}(x_{2},a)=\pi.$
By symmetry, $F(a)$ is constant along the great circle orthogonal to the axis through $x_{1}$ and $x_{2}$. Hence, the intrinsic Fréchet mean is not unique: the Fréchet mean set is precisely this equatorial great circle.
For example, choosing
$a^{*}=(0,1,0)\quad\text{or}\quad a^{*}=(0,-1,0),$
we have
$d_{G}(x_{i},a^{*})=\frac{\pi}{2},\quad i=1,2,$
and therefore
$V_{1}(P)=\frac{1}{2}\left(\frac{\pi}{2}\right)^{2}+\frac{1}{2}\left(\frac{\pi}{2}\right)^{2}=\frac{\pi^{2}}{4}.$
Extrinsic centroid. The extrinsic centroid (see Section 1.13) is obtained by first computing the Euclidean mean
$m=\tfrac{1}{2}(x_{1}+x_{2})=0\in\mathbb{R}^{3}.$
Since $m=0$, normalization onto the sphere is not possible, and hence the extrinsic centroid is undefined in this case.
Two-means case. For $n=2$, choosing the codebook $\alpha=\{x_{1},x_{2}\}$ yields zero distortion, and thus
$V_{2}(P)=0.$
Remark. This example illustrates that for antipodally symmetric distributions on the sphere, intrinsic Fréchet means may be non-unique, while extrinsic centroids may fail to exist. Such behavior has no Euclidean analogue and highlights the geometric nature of quantization on spherical manifolds.
@@ 11. Further Remarks and Extensions
  • Nonuniform densities. For nonuniform distributions on an arc or circle, the Voronoi cells are no longer equal in length; their boundaries shift to equalize weighted distortion. The analysis involves solving $f(x)(x-a_{i})$ equilibrium equations along geodesics.